OpenStudy (anonymous):

factor this trinomial; 2x^2-7x-30.

OpenStudy (anonymous):

please show the steps. thxx

OpenStudy (anonymous):

so gcf is 1. then trial and error?

OpenStudy (saifoo.khan):

Sure, 2x^2-7x-30. 2x^2 - 12x +5x -30 = 0 2x(x-6) +5(x-6) = 0 (2x+5)(x-6)= 0

OpenStudy (anonymous):

so to solve this equation, as opposed to factoring, is the result different?